How many grams of water in a cup that requires 4,500J to heat from 20C to 40C?
1
Expert's answer
2020-05-15T12:10:51-0400

Solution:

For Q = m × C × ΔT,

Q = amount of heat (J)

m = mass (grams)

C = specific heat (J/g°C)

ΔT = change in temperature (°C)


Q = 4500 J

C(H2O) = 4.186 J/g°C

ΔT = Tfinal - Tinitial = (40 - 20)°C = 20°C


Thus,

m = Q / (C × ΔT)

m = (4500 J) / (4.186 J/g°C × 20°C) = 53.75 g

m(H2O) = 53.75 g


Answer: 53.75 grams of water.